OpenStudy (anonymous):

What's the median for the following set of data. 4, 6, 2, 9, 6, 1

OpenStudy (anonymous):

Median is the middle. rearrange them in order from least to greatest. 1,2,4,6,6,9 in this case there is no median within the numbers but the median would be 5.

OpenStudy (anonymous):

Median is 5. Arrange the numbers orderly to get; 1, 2, 4, 6, 6, 9 and crossing one from each side you get 4 and 6 left in the middle 4+6 = 10/2 = 5.

OpenStudy (anonymous):

So it's 5.5

OpenStudy (anonymous):

no just 5

OpenStudy (anonymous):

Ok